OpenStudy (jessiegonzales):

medal!!!! 1. To which set of numbers does the number belong? -55 A. natural numbers B. whole numbers C. integers D. irrational numbers

OpenStudy (jessiegonzales):

C?

OpenStudy (jessiegonzales):

@nincompoop

OpenStudy (nincompoop):

whole and natural, \(\sf N\), are \(0 \rightarrow \infty \) and \(1 \rightarrow \infty \) integers, \(\sf Z^+~and ~Z^-\), are \(-\infty \leftarrow 0 \rightarrow + \infty\) irrational, are numbers that contains decimals with no repeating pattern (non-terminal)

OpenStudy (jessiegonzales):

thanks!! Answer was C